OpenStudy (anonymous):

Passing through (3,4) with x intercept 4 what is the? point slope and slope intercept form?

OpenStudy (anonymous):

if this is a linear equation which I am assuming it is. the slope of the equation is rise over run. so then the rise is actually negative or -3 and the run is 1 so the slope is -3x to find the equation for now all you have to do is use y-y1=m(x-x1) now you know that m=-3 and you have both y1 and x1 given to you by the point. I really don't remember the difference between point slope and slope intercept forms, but I think you should be able to figure it out with this.